If f(x) = x2 − 8x + 16, find f(−a), f(a − 4), and f(a + h).
f(−a) =
f(a − 4) =
f(a + h) =

Answer by Theo(13342)

f(x) = x^2 - 8x + 16

f(-a) = (-a)^2 - 8(-a) + 16

f(a-4) = (a-4)^2 - 8(a-4) + 16

f(a + h) = (a + h)^2 - 8(a + h) + 16

x is the argument for f(x).

-a is the argument for f(-a).

a-4 is the argument for f(a-4).

a+h is the argument for f(a + h).

in functional notation, the formula is working on the argument.

when you see f(x) = x^2 - 8x + 16, the argument is x and that's what the formula is working on.

when you see f(-a) = (-a)^2 - 8(-a) + 16, the argument is (-a) and that what the formula is working on.

the formula remains the same.
the only thing different is the argument that it's working on.

an alternate way of telling you this:

start with f(x) = x^2 - 8x + 16
you want to find the value of the function when x = 9.
your new function becomes:
f(9) = 9^2 - 8*9 + 16

this is no different.

for f(-a), you want to find the value of the function when x = -a.

for f(a-4), you want to find the value of the function when x = (a-4).

for f(a+h), you want to find the value of the function when x = (a+h).

the function itself is the rules that operate on the argument provided.

the rules of this function are:

take the argument and then square it.
then take the argument and multiply it by 8 and subtract it.
then take 16 and add it.

the argument is whatever you want it to be.
it starts off with x.
when you are supplied with a different argument, you just replace x with that argument.
the function remains the same.